OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. Type your answers in the blanks. |4x+3|=9+2x x = _____ or _____

OpenStudy (anonymous):

first you need to do this: 4x+3=9+2x & 4x+3=-9+2x now you solve each equation for x :)